OpenStudy (anonymous):

Find the 11th term of the arithmetic sequence. a3 = 20.5 and a8 = 13, a11 = ______.

OpenStudy (cornzyblack):

ok uhmmm first we gata use the formula which i believe is a + (n - 1)d so if a3 = 20.5 then, a3 = a + (n - 1)d n= 3 20.5 = a + (3 - 1)d ----- 1 20.5 = a + 2d ----- 1 then, a8 = a + (n - 1)d n= 8 13 = a + (8 - 1)d -------- 2 13 = a + 7d --------- 2 So we are meant to find what a11 is n = 11, a =? a11 = a + (11-1)d a11 = a - 10d -------3 Well we have a simultaneous eqn already So, 20.5 = a + 2d 13 = a + 7d now lets solve for a in eqn 1 a = 20.5 - 2d -------- 4 substitute in eqn 2 so, 13 = (20.5 - 2d) + 7d 13 = 20.5 + 5d 5d = 13 - 20.5 5d = -7.5 d = -7.5/5 So sub in eqn 1 or 2 to get a so using eqn 2 a = 13 - 7d a = 13 - 7(-7.5/5) a = 23.5 so we have a and d hence a11 = a +(n - 1)d if a = 23.5 , n = 11 then a11 = 23.5 + (11 -1)(-7.5/ 5) a11 = 23.5 + (10)(-1.5)

OpenStudy (cornzyblack):

a 11 = 8.5

OpenStudy (lilg132):

a + 2d = 20.5 a + 8d = 13 -6d = 7.5 -d = 1.25 d = -1.25 a - 2.5 = 20.5 a = 20.5 + 2.5 a = 23 a + d = 23 - 1.25 = 21.75 is a2 23 +2(-1.25) = 20.5 is a3 23 + 8(-1.25) = 13 is a8 23 + 11(-1.25) = 9.25 is a11 Please remember to click good answer

OpenStudy (lilg132):

your answer is wrong cornzyblack

OpenStudy (lilg132):

and your formula is also incorrect

OpenStudy (cornzyblack):

its called openstudy

OpenStudy (lilg132):

when did you realise?

OpenStudy (cornzyblack):

mr. lig i tink u gat beef with me u want me to show u a flaw in your work, check line 2, u'll see u did not subtract to get a + 7d = a8, so back off. almost everytym i come here some1 must piss me off

OpenStudy (lilg132):

a + 7d = 13 your write i made a typo but atleast i can admit my mistake admit your wrong and your answer is also wrong as well as your method and we have nothing further to talk about lol

OpenStudy (cornzyblack):

am definitely sure am not wrong

OpenStudy (lilg132):

lol are you saying im wrong then?

OpenStudy (anonymous):

Well, Idk if cornzy is right or not, but lilg, your answer was incorrect. I used yours and got that question wrong.

OpenStudy (lilg132):

yes i'm working it out now because of my typo i have to do it again

OpenStudy (cornzyblack):

see lilg if am wrong i would like someone to tell me not u from u, k. Beause you just pointed tha finger at me when u did rubbish

OpenStudy (lilg132):

well done cornzyblack you got it right

OpenStudy (lilg132):

a11 = 8.5